# Topology and Diffeology via Metric-like Functions

### v2

@@ 1. Introduction
The space $\Psi_{d}(\mathbb{R}^{N})$ of all closed $d$-dimensional submanifolds in $\mathbb{R}^{N}$ has been studied from various perspectives. Among these, Galatius–Randal-Williams introduced in [8] a specific topology on $\Psi_{d}(\mathbb{R}^{N})$. Later, in Schommer-Pries [11], it was shown that this topology can be reinterpreted as the $D$-topology of a certain diffeology on $\Psi_{d}(\mathbb{R}^{N})$, providing a new perspective for studying this space.
Diffeological spaces, introduced in [13] by Souriau in the 1980s, extend the notion of smooth manifolds to a broader class of spaces, including singular spaces and infinite-dimensional objects. A diffeology naturally induce a topology, the $D$-topology, which is compatible with the smooth structure and provides a way to extend smooth structures while respecting the underlying topology. This makes them particularly useful for studying spaces like $\Psi_{d}(\mathbb{R}^{N})$, where standard manifold structures are not directly applicable.
Fundamental properties of this topology, including its metrizability, were first established by Bökstedt-Madsen in [3]. In [3], a family of functions $\{d_{r}:\Psi_{d}(\mathbb{R}^{N})\times\Psi_{d}(\mathbb{R}^{N})\to\mathbb{R}_{\geq 0}\cup\{\infty\}\}_{r>0}$, indexed by $r>0$, was introduced to analyze the topology of $\Psi_{d}(\mathbb{R}^{N})$. These functions satisfy properties which resemble the axioms of a metric:
Proposition 1.1 ([3], Lemma 4.2).
@@
Furthermore, in [3], it was shown that these functions define a topology on $\Psi_{d}(\mathbb{R}^{N})$, which coincides with the topology defined in [8].
Proposition 1.2 ([3], Lemma 4.3, Theorem 2.1).
@@
Then, the collection of sets $\mathcal{U}_{r,\varepsilon}(M)$ forms a basis for a topology on $\Psi_{d}(\mathbb{R}^{N})$, which coincides with the topology introduced in [8].
Motivated by these developments, this paper introduces an abstract framework for spaces equipped with families of metric-like functions $\{d_{r}\}_{r>0}$, satisfying axioms similar to those in [3]. We show that for such spaces, topology, uniformity, and diffeology can be defined in a unified manner. This approach not only generalizes the structures found on $\Psi_{d}(\mathbb{R}^{N})$ but also provides new insights into other spaces.
One example we focus on in this paper is the tiling space, which is well studied and described in [10] by Sadun as a metric space representing a space of tilings of $\mathbb{R}^{d}$. In [1] by Alatorre–Rodríguez-Guzmán, a diffeology was defined on tiling spaces, and its basic properties were studied. We show that the topology on tiling spaces, as defined in [10], can naturally be reinterpreted as arising from a family of $\{d_{r}\}_{r>0}$. Moreover, we show that the diffeology defined using this family $\{d_{r}\}_{r>0}$ coincides with the one introduced in [1]. These results form the main theorem of this paper.
Additionally, we show that the axioms for $\{d_{r}\}_{r>0}$ introduced in this paper appear in various other contexts, including the compact-open topology and the space of all graphs in $\mathbb{R}^{N}$. In the case of the space of graphs, as studied by Galatius in [7], a slight relaxation of the axioms—specifically, a further weakening of the triangle inequality—allows us to incorporate these examples naturally. Thus, this paper provides a unified framework for studying spaces defined by families of metric-like functions, offering a new perspective on their topology, uniformity, and diffeology. This work represents a first step toward a broader understanding of such spaces, with potential applications beyond the specific cases considered here.
We also show that a family $\{d_{r}\}_{r>0}$ gives rise to a natural pseudo-metric that is not necessarily symmetric on the space. In particular, in the case of submanifolds, this provides a simpler way to recover the result of [3], Theorem 4.4.
@@ 2. Topology defined by $d_{r}$
We first introduce the following axioms, inspired by the properties stated in Proposition 1.1, which were originally established in [3].
@@
Let $X$ be a set, and $\{d_{r}\colon X\times X\to\mathbb{R}_{\geq 0}\cup\{\infty\}\}_{r>0}$ be a family of functions indexed by $r>0$. We assume that the following axioms hold for any three elements $x$, $y$, $z\in X$:
@@
The following proposition and its proof are identical to those given in [3].
Proposition 2.4 ([3], Lemma 4.3).
@@
$d_{r}(x,\cdot)\colon X\to\mathbb{R};y\mapsto d_{r}(x,y)$
@@
Let $U\subset\mathbb{R}$ be any open interval of the form $(-\infty,b)$, and let $y\in d_{r}(x,\cdot)^{-1}(U)=\mathcal{U}_{r,b}(x)$ be any element. By the Proposition 2.4, there exist $r^{\prime}>0$ and $\varepsilon>0$ such that
@@
This implies that $d_{r}(x,\cdot)\colon X\to\mathbb{R}$ is upper semi-continuous.
@@
Let $K\subset T$ be an arbitrary compact subset, and let $U\subset S$ be an arbitrary open subset. Consider an element $f\in W(K,U)=\{g\colon T\to S\mid g(K)\subset U\}$. Since $K$ is compact, we can choose $r>0$ such that $K\subset{r}D$. For each $x\in K$, there exists $\varepsilon_{x}>0$ such that $B_{\varepsilon_{x}}{(f(x))}\subset U$. By compactness of $K$, we can choose a finite set $\{x_{1},\dots,x_{n}\}\subset K$ such that
@@
For the reverse inclusion, let $r>0$ and $\varepsilon>0$ be arbitrary, and consider $f\in C^{0}(T,S)$. For each $x\in{r}D$, there exists $r_{x}>0$ such that $B_{2r_{x}}{(x)}\subset f^{-1}(B_{\varepsilon/2}{(f(x))})$. Since ${r}D$ is compact (by the assumption that $T$ is proper), there exists a finite set $\{x_{1},\dots,x_{n}\}\subset{r}D$ such that
@@
$\Psi_{d}(\mathbb{R}^{N})=\{W^{d}\subset\mathbb{R}^{N}\mid\mbox{$W^{d}$ is a closed submanifold of $\mathbb{R}^{N}$ such that $\partial W=\emptyset$.}\}$
@@
$\mathcal{P}_{r}(V,W)=\left\{\phi\colon U\to\phi(U)\;\middle|\begin{array}[]{l}\mbox{$U\subset V$ and $\phi(U)\subset W$ are both open,}\\ \mbox{$\phi\colon U\to\phi(U)$ is a diffeomorphism,}\\ \mbox{$V\cap{r}D\subset U$,\; $W\cap{r}D\subset\phi(U)$.}\end{array}\right\}.$
@@
for each $r>0$. If $\mathcal{P}_{r}(V,W)=\emptyset$, we set $d_{r}(V,W)=\infty$. This family $\{d_{r}\}_{r>0}$ satisfies the axioms in Definition 2.1. The topology induced by this family $\{d_{r}\}_{r>0}$ coincides with the one defined in [8].
@@
This result was proved in [3], so we omit the proof here. ∎
@@ 3. Tiling spaces and $d_{r}$
In this section, we recall the definition of tiling spaces and discuss the relationship between the tiling spaces and the family $\{d_{r}\}_{r>0}$. For a more detailed treatment of tiling spaces, we refer the reader to the textbook [10] and the paper [2] by Anderson-Putnam.
Definition 3.1 ([10], [2], [1]).
@@
  • For each $t_{1},t_{2}\in T$ such that $t_{1}\neq t_{2}$, $\operatorname{Int}{t_{1}}\cap\operatorname{Int}{t_{2}}\neq\emptyset$.
@@
$d(T,T^{\prime})=\inf\left(\left\{\varepsilon>0\;\middle|\begin{array}[]{l}\mbox{There exist $u,v\in\mathbb{R}^{d}$ such that $|u|<\varepsilon$, $|v|<\varepsilon$,}\\ \mbox{and $(T+u)\cap B_{1/\varepsilon}{(0)}=(T^{\prime}+v)\cap B_{1/\varepsilon}{(0)}$}\end{array}\right\}\cup\{1/\sqrt{2}\}\right).$
This $d$ is easily verified to be a metric on $O{(T)}$. We define the tiling space $\Omega_{T}$ of $T$ as the metric completion of $O{(T)}$ with respect to $d$.
@@
If $T$ is tiling with $\mathscr{P}=\{\tau_{1},\dots,\tau_{n}\}$ as its set of prototiles, then any $T^{\prime}\in\Omega_{T}$ can be regarded as a tiling whose set of prototiles is a subset of $\mathscr{P}$. This is because any limit of a Cauchy sequence in $O{(T)}$ corresponds to a tiling in which every patch appearing in $T^{\prime}$ can also be found in $T$. Here, a patch refers to a finite and connected union of tiles.
@@
$d_{r}(T,T^{\prime})=\inf\left\{\varepsilon>0\;\middle|\begin{array}[]{l}\mbox{There exists $u\in\mathbb{R}^{d}$ such that $|u|<\varepsilon$,}\\ \mbox{and $(T+u)\cap{r}D=T^{\prime}\cap{r}D$}\end{array}\right\}.$
@@
$\bigcup\{t\in(T_{1}+u)\cup T_{2}\mid t\cap{r}D\neq\emptyset\}$
is closed, we can separate it from ${r}D$ using the normality of $\mathbb{R}^{d}$. Therefore, there exists $\delta>0$ such that $(T_{1}+u)\cap{(r+\delta)}D=T_{2}\cap{(r+\delta)}D$. This implies $d_{r+\delta}(T_{1},T_{2})<\varepsilon$.
@@
It suffices to prove that $d_{r_{1}}(T_{1},T_{3})\leq d_{1}+d_{2}$. First, there exist $u,v\in\mathbb{R}^{d}$ such that $|u|<d_{1}$, $|v|<d_{2}$, and $(T_{1}+u)\cap{r_{12}}D=T_{2}\cap{r_{12}}D$, $(T_{2}+v)\cap{r_{123}}D=T_{2}\cap{r_{123}}D$. Using these $u$ and $v$, we now verify the following identity to prove $d_{r_{1}}(T_{1},T_{3})\leq d_{1}+d_{2}$:
@@
Let $T_{1}$ be a tiling of $\Omega_{T}$. For any $r>0$ and $\varepsilon>0$, we need to find a number $\varepsilon^{\prime}>0$ such that $B_{\varepsilon^{\prime}}{(T_{1})}\subset\mathcal{U}_{r,\varepsilon}(T_{1})$, where $B_{\varepsilon^{\prime}}{(T_{1})}=\{T^{\prime}\in\Omega_{T}\mid d(T_{1},T^{\prime})<\varepsilon^{\prime}\}$. Let $\mathscr{P}$ be the set of all prototiles of $T$. By Remark 3.3, any tiling in $\Omega_{T}$ consists of tiles that are translated copies of the prototiles in $\mathscr{P}$. We define $\varepsilon^{\prime}>0$ to be a positive number satisfying
@@ 4. Uniformity defined by $d_{r}$
The family $\{d_{r}\}_{r>0}$ induces not only a topology but also a quasi-uniformity. Moreover, an even weaker version of the triangle inequality suffices to define the quasi-uniformity. In particular, this weaker triangle inequality is satisfied in the space of graph embeddings $\Phi{(\mathbb{R}^{N})}$, as defined in [7].
Before proceeding, we provide a brief introduction to quasi-uniform spaces, which generalize uniform spaces by not necessarily requiring symmetry. For a more detailed treatment of uniform spaces, see [4], Chapter II by Bourbaki, and for quasi-uniform spaces, see [6] by Fletcher-Lindgren.
Definition 4.1 ([4], Chapter II, [6], Chapter 1).
@@
Proposition 4.2 ([4], Chapter II, §1.2, Proposition 1).
@@
Although [4] states this proposition for uniform spaces, the proof does not rely on the symmetry of the entourages. Thus, the statement also holds for quasi-uniform spaces.
@@
As an example, we show how the family $\{d_{r}\}_{r>0}$, satisfying the weaker triangle inequality, can describe the $C^{0}$-topology on the space $\Phi{(\mathbb{R}^{N})}$ of all graphs in $\mathbb{R}^{N}$, as defined in [7]. For consistency, we follow the definitions and notation in [7] without restating them here.
@@
Let $\Phi{(\mathbb{R}^{N})}$ be the space of all graphs $G$ defined in [7]. For each $r>0$ and $G,H\in\Phi{(\mathbb{R}^{N})}$, we define
@@
Moreover, for each $\varphi\in\mathcal{P}_{r}(V,W)$, we define $s(\varphi,r)=(H\cap{r}D)\cup\varphi^{-1}(G\cap{r}D)$, and
@@
for each $r>0$. If $\mathcal{P}_{r}(G,H)=\emptyset$, we set $d_{r}(G,H)=\infty$. This family $\{d_{r}\}_{r>0}$ satisfies the weaker $\{d_{r}\}$-axioms in Definition 4.4. The topology induced by $\{d_{r}\}_{r>0}$ coincides with the $C^{0}$-topology defined in [7], Definition 2.5.
@@
In this example, the family $\{d_{r}\}_{r>0}$ may not satisfy the symmetry condition $d_{r}(G,H)=d_{r}(H,G)$. Moreover, note that the domain of the triple $\varphi$ in $\mathcal{P}_{r}(G,H)$ is an open subset of $H$, not $G$. Additionally, [7] defines a $C^{1}$-topology for the space $\Phi{(\mathbb{R}^{N})}$. However, regardless of how one defines the family $\{d_{r}\}_{r>0}$, it seems difficult to recover this topology from it.
@@
Although the proof is largely similar to Lemma 4.2, in [3], some conditions have been modified. Therefore, we provide the proof here for completeness.
@@
Let $r>0$, $G,H\in\Psi_{d}(\mathbb{R}^{N})$, and $\varepsilon>0$ be arbitrary satisfying $d_{r}(G,H)<\varepsilon$. Then, there exists a map $\varphi\in\mathcal{P}_{r}(G,H)$ with $\varphi\colon U\to\phi(U)$ satisfying $|p-\varphi(p)|<\varepsilon$ for all $p\in s(\varphi,r)$. Since $U$ is open in $H$, the set
@@
Next, we prove the weak triangle inequality.
Let $G_{1},G_{2},G_{3}\in\Phi(\mathbb{R}^{N})$, and $r_{1},r_{2},r_{3}>0$ be arbitrary. Suppose that
@@
There exist two maps $(\varphi_{1}\colon U_{1}\to\varphi_{1}(U_{1}))\in\mathcal{P}_{r_{123}}(G_{1},G_{2})$ and $(\varphi_{2}\colon U_{2}\to\varphi(U_{2}))\in\mathcal{P}_{r_{123}}(G_{2},G_{3})$ such that
@@
For any $p_{3}\in s(\varphi_{2},r_{13})$, we have either $p_{3}\in G_{3}\cap{r_{13}}D$ or $\varphi_{2}(p_{3})\in G_{2}\cap{r_{13}}D\subset G_{2}\cap{r_{123}}D$. Even if $p_{3}\in G_{3}\cap{r_{13}}D$, the inequality $|p_{3}-\varphi_{2}(p_{3})|<d_{2}<r_{2}$ implies that $\varphi_{2}(p_{3})\in G_{2}\cap{r_{123}}D$. Therefore, for any $p_{3}\in s(\varphi_{1},r_{13})$, we obtain $\varphi_{2}(p_{3})\in G_{2}\cap{r_{123}}D$, which allows us to restrict the domain of $\varphi_{2}$ to a smaller neighborhood $U$ that contains $s(\varphi_{2},r_{13})$ and on which the composition $\varphi_{1}\circ\varphi_{2}$ is well-defined.
@@
Finally, we prove that the topology induced by $\{d_{r}\}_{r>0}$ coincides with the $C^{0}$-topology defined in [7], Definition 2.5. First, we show that for any $G\in\Phi{(\mathbb{R}^{N})}$, any compact set $K\subset\mathbb{R}^{N}$, and any $\varepsilon>0$, there exist $r^{\prime}>0$ and $\varepsilon^{\prime}>0$ such that

### v3

@@ 1. Introduction
The space $\Psi_{d}(\mathbb{R}^{N})$ of all closed $d$-dimensional submanifolds in $\mathbb{R}^{N}$ has been studied from various perspectives. Among these, Galatius–Randal-Williams introduced in [MR2653727] a specific topology on $\Psi_{d}(\mathbb{R}^{N})$, based on ideas that first appeared implicitly in the study of the cobordism category by Galatius-Madsen-Tillmann-Weiss [MR2506750] and Galatius [MR2784914]. Later, in Schommer-Pries [MR4821355], it was shown that this topology can be reinterpreted as the $D$-topology of a certain diffeology on $\Psi_{d}(\mathbb{R}^{N})$, providing a new perspective for studying this space.
Diffeological spaces, introduced in [MR0607688] by Souriau in the 1980s, extend the notion of smooth manifolds to a broader class of spaces, including singular spaces and infinite-dimensional objects. Any diffeology naturally induces a topology, the $D$-topology, which is compatible with the smooth structure and provides a way to extend smooth structures while respecting the underlying topology. This makes them particularly useful for studying spaces like $\Psi_{d}(\mathbb{R}^{N})$, where standard manifold structures are not directly applicable.
Fundamental properties of this topology, including its metrizability, were first established by Bökstedt-Madsen in [MR3243393]. In [MR3243393], a family of functions $\{d_{r}\colon\Psi_{d}(\mathbb{R}^{N})\times\Psi_{d}(\mathbb{R}^{N})\rightarrow\mathbb{R}_{\geq 0}\cup\{\infty\}\}_{r>0}$, indexed by $r>0$, was introduced to analyze the topology of $\Psi_{d}(\mathbb{R}^{N})$. These functions satisfy properties which resemble the axioms of a metric:
Proposition 1.1 ([MR3243393], Lemma 4.2).
@@
Furthermore, in [MR3243393], it was shown that these functions define a topology on $\Psi_{d}(\mathbb{R}^{N})$, which coincides with the topology defined in [MR2653727].
Proposition 1.2 ([MR3243393], Lemma 4.3, Theorem 2.1).
@@
Then, the collection of sets $\mathcal{U}_{r,\varepsilon}(M)$ forms a basis for a topology on $\Psi_{d}(\mathbb{R}^{N})$, which coincides with the topology introduced in [MR2653727].
Motivated by these developments, this paper introduces an abstract framework for spaces equipped with families of metric-like functions $\{d_{r}\}_{r>0}$, satisfying axioms similar to those in [MR3243393]. We show that for such spaces, topology, uniformity, and diffeology can be defined in a unified manner. This approach not only generalizes the structures found on $\Psi_{d}(\mathbb{R}^{N})$ but also provides new insights into other spaces.
One example we focus on in this paper is the tiling space, which is well studied and described in [MR2446623] by Sadun as a metric space representing a space of tilings of $\mathbb{R}^{d}$. In [MR4882913] by Alatorre–Rodríguez-Guzmán, a diffeology was defined on tiling spaces, and its basic properties were studied. We show that the topology on tiling spaces, as defined in [MR2446623], can naturally be reinterpreted as arising from a family of $\{d_{r}\}_{r>0}$. Moreover, we show that the diffeology defined using this family $\{d_{r}\}_{r>0}$ coincides with the one introduced in [MR4882913]. These results form the main theorem of this paper.
Additionally, we show that the axioms for $\{d_{r}\}_{r>0}$ introduced in this paper appear in various other contexts, including the compact-open topology and the space of all graphs in $\mathbb{R}^{N}$. In the case of the space of graphs, as studied by Galatius in [MR2784914], a slight relaxation of the axioms—specifically, a further weakening of the triangle inequality—allows us to incorporate these examples naturally. Thus, this paper provides a unified framework for studying spaces defined by families of metric-like functions, offering a new perspective on their topology, uniformity, and diffeology. This work represents a first step toward a broader understanding of such spaces, with potential applications beyond the specific cases considered here.
We also show that a family $\{d_{r}\}_{r>0}$ gives rise to a natural pseudo-metric that is not necessarily symmetric on the space. In particular, in the case of submanifolds, this provides a simpler way to recover the result of [MR3243393], Theorem 4.4. Cantero [Cantero] also provided an explicit metric that induces the same topology on submanifold spaces. Our approach takes a different angle by using an axiomatic formulation, which can be naturally applied to other types of spaces as well.
@@ 2. Topology defined by $d_{r}$
We first introduce the following axioms, inspired by the properties stated in Proposition 1.1, which were originally established in [MR3243393].
@@
Let $X$ be a set, and $\{d_{r}\colon X\times X\rightarrow\mathbb{R}_{\geq 0}\cup\{\infty\}\}_{r>0}$ be a family of functions indexed by $r>0$. We assume that the following axioms hold for any three elements $x$, $y$, $z\in X$:
@@
The following proposition and its proof are identical to those given in [MR3243393].
Proposition 2.4 ([MR3243393], Lemma 4.3).
@@
$d_{r}(x,\cdot)\colon X\rightarrow\mathbb{R}_{\geq 0}\cup\{\infty\},\quad y\mapstochar\rightarrow d_{r}(x,y)$
@@
Let $b\in\mathbb{R}$, and let $y\in d_{r}(x,\cdot)^{-1}((-\infty,b))=\mathcal{U}_{r,b}(x)$ be any element. By Proposition 2.4, there exist $r^{\prime}>0$ and $\varepsilon>0$ such that
@@
This implies that $d_{r}(x,\cdot)\colon X\rightarrow\mathbb{R}_{\geq 0}\cup\{\infty\}$ is upper semi-continuous. ∎
@@
Let $K\subset T$ be an arbitrary compact subset, and let $U\subset S$ be an arbitrary open subset. Consider an element $f\in W(K,U)=\{g\colon T\rightarrow S\mid g(K)\subset U\}$. Since $K$ is compact, we can choose $r>0$ such that $K\subset{r}D$. For each $x\in K$, there exists $\varepsilon_{x}>0$ such that $B_{\varepsilon_{x}}{(f(x))}\subset U$. By compactness of $K$, we can choose a finite set $\{x_{1},\dots,x_{n}\}\subset K$ such that
@@
For the reverse inclusion, let $r>0$ and $\varepsilon>0$ be arbitrary, and consider $f\in C^{0}(T,S)$. For each $x\in{r}D$, there exists $r_{x}>0$ such that $B_{2r_{x}}{(x)}\subset f^{-1}(B_{\varepsilon/2}{(f(x))})$. Since rD{r}D is compact (by the assumption that $T$ is proper), there exists a finite set $\{x_{1},\dots,x_{n}\}\subset{r}D$ such that
@@
[…]
@@
[…]
@@
for each $r>0$. If $\mathcal{P}_{r}(V,W)=\emptyset$, we set $d_{r}(V,W)=\infty$. This family $\{d_{r}\}_{r>0}$ satisfies the axioms in Definition 2.1. The topology induced by this family $\{d_{r}\}_{r>0}$ coincides with the one defined in [MR2653727].
@@
This result was proved in [MR3243393], so we omit the proof here. ∎
@@ 3. Tiling spaces and $d_{r}$
In this section, we recall the definition of tiling spaces and discuss the relationship between the tiling spaces and the family $\{d_{r}\}_{r>0}$. For a more detailed treatment of tiling spaces, we refer the reader to the textbook [MR2446623] and the paper [MR1631708] by Anderson-Putnam.
Definition 3.1 ([MR2446623], [MR1631708], [MR4882913]).
@@
  • For each $t_{1},t_{2}\in T$ such that $t_{1}\neq t_{2}$, $\operatorname{Int}{t_{1}}\cap\operatorname{Int}{t_{2}}=\emptyset$.
@@
[…]
This $d$ is easily verified to be a metric on O(T)O{(T)}. We define the tiling space $\Omega_{T}$ of $T$ as the metric completion of O(T)O{(T)} with respect to $d$.
@@
If $T$ is a tiling with $\mathscr{P}=\{\tau_{1},\dots,\tau_{n}\}$ as its set of prototiles, then any $T^{\prime}\in\Omega_{T}$ can be regarded as a tiling whose set of prototiles is a subset of $\mathscr{P}$. This is because any limit of a Cauchy sequence in O(T)O{(T)} corresponds to a tiling in which every patch appearing in $T^{\prime}$ can also be found in $T$. Here, a patch refers to a finite and connected union of tiles.
@@
$d_{r}(T_{1},T_{2})=\inf\left\{\varepsilon>0\;\middle|\begin{array}[]{l}\mbox{There exists $u\in\mathbb{R}^{d}$ such that $|u|<\varepsilon$,}\\ \mbox{and $(T_{1}+u)\cap{r}D=T_{2}\cap{r}D$}\end{array}\right\}.$
@@
$\bigcup\{t\in(T_{1}+u)\cup T_{2}\mid t\cap{r}D=\emptyset\}$
is closed, we can separate it from rD{r}D using the normality of $\mathbb{R}^{d}$. Therefore, there exists $\delta>0$ such that $(T_{1}+u)\cap{(r+\delta)}D=T_{2}\cap{(r+\delta)}D$. This implies $d_{r+\delta}(T_{1},T_{2})<\varepsilon$.
@@
It suffices to prove that $d_{r_{1}}(T_{1},T_{3})\leq d_{1}+d_{2}$. First, there exist $u,v\in\mathbb{R}^{d}$ such that $|u|<d_{1}$, $|v|<d_{2}$, and $(T_{1}+u)\cap{r_{12}}D=T_{2}\cap{r_{12}}D$, $(T_{2}+v)\cap{r_{123}}D=T_{3}\cap{r_{123}}D$. Using these $u$ and $v$, we now verify the following identity to prove $d_{r_{1}}(T_{1},T_{3})\leq d_{1}+d_{2}$:
@@
Let $T_{1}$ be a tiling in $\Omega_{T}$. For any $r>0$ and $\varepsilon>0$, we need to find a number $\varepsilon^{\prime}>0$ such that $B_{\varepsilon^{\prime}}{(T_{1})}\subset\mathcal{U}_{r,\varepsilon}(T_{1})$, where $B_{\varepsilon^{\prime}}{(T_{1})}=\{T^{\prime}\in\Omega_{T}\mid d(T_{1},T^{\prime})<\varepsilon^{\prime}\}$. Let $\mathscr{P}$ be the set of all prototiles of $T$. By Remark 3.3, any tiling in $\Omega_{T}$ consists of tiles that are translated copies of the prototiles in $\mathscr{P}$. We define $\varepsilon^{\prime}>0$ to be a positive number satisfying
@@ 4. Uniformity defined by $d_{r}$
The family $\{d_{r}\}_{r>0}$ induces not only a topology but also a quasi-uniformity. Moreover, an even weaker version of the triangle inequality suffices to define the quasi-uniformity. In particular, this weaker triangle inequality is satisfied in the space of graph embeddings $\Phi{(\mathbb{R}^{N})}$, as defined in [MR2784914].
Before proceeding, we provide a brief introduction to quasi-uniform spaces, which generalize uniform spaces by not necessarily requiring symmetry. For a more detailed treatment of uniform spaces, see [MR1726779], Chapter II by Bourbaki, and for quasi-uniform spaces, see [MR660063] by Fletcher-Lindgren.
Definition 4.1 ([MR1726779], Chapter II, [MR660063], Chapter 1).
@@
Proposition 4.2 ([MR1726779], Chapter II, §1.2, Proposition 1).
@@
Although [MR1726779] states this proposition for uniform spaces, the proof does not rely on the symmetry of the entourages. Thus, the statement also holds for quasi-uniform spaces.
@@
As an example, we show how the family $\{d_{r}\}_{r>0}$, satisfying the weaker triangle inequality, can describe the $C^{0}$-topology on the space $\Phi{(\mathbb{R}^{N})}$ of all graphs in $\mathbb{R}^{N}$, as defined in [MR2784914]. For consistency, we follow the definitions and notation in [MR2784914] without restating them here.
@@
Let $\Phi{(\mathbb{R}^{N})}$ be the space of all graphs $G$ defined in [MR2784914]. For each $r>0$ and $G,H\in\Phi{(\mathbb{R}^{N})}$, we define
@@
Moreover, for each $\varphi\in\mathcal{P}_{r}(G,H)$, we define $s(\varphi,r)=(H\cap{r}D)\cup\varphi^{-1}(G\cap{r}D)$, and
@@
for each $r>0$. If $\mathcal{P}_{r}(G,H)=\emptyset$, we set $d_{r}(G,H)=\infty$. This family $\{d_{r}\}_{r>0}$ satisfies the weaker $\{d_{r}\}$-axioms in Definition 4.4. The topology induced by $\{d_{r}\}_{r>0}$ coincides with the $C^{0}$-topology defined in [MR2784914], Definition 2.5.
@@
In this example, the family $\{d_{r}\}_{r>0}$ may not satisfy the symmetry condition $d_{r}(G,H)=d_{r}(H,G)$. Moreover, note that the domain of the triple $\varphi$ in $\mathcal{P}_{r}(G,H)$ is an open subset of $H$, not $G$. Additionally, [MR2784914] defines a $C^{1}$-topology for the space $\Phi{(\mathbb{R}^{N})}$. However, regardless of how one defines the family $\{d_{r}\}_{r>0}$, it seems difficult to recover this topology from it.
@@
Although the proof is largely similar to Lemma 4.2, in [MR3243393], some conditions have been modified. Therefore, we provide the proof here for completeness.
@@
Let $r>0$, $G,H\in\Phi{(\mathbb{R}^{N})}$, and $\varepsilon>0$ be arbitrary satisfying $d_{r}(G,H)<\varepsilon$. Then, there exists a map $\varphi\in\mathcal{P}_{r}(G,H)$ with $\varphi\colon U\rightarrow\varphi(U)$ satisfying $|p-\varphi(p)|<\varepsilon$ for all $p\in s(\varphi,r)$. Since $U$ is open in $H$, the set
@@
Next, we prove the weaker triangle inequality. Let $G_{1},G_{2},G_{3}\in\Phi(\mathbb{R}^{N})$, and $r_{1},r_{2},r_{3}>0$ be arbitrary. Suppose that
@@
There exist two maps $(\varphi_{1}\colon U_{1}\rightarrow\varphi_{1}(U_{1}))\in\mathcal{P}_{r_{123}}(G_{1},G_{2})$ and $(\varphi_{2}\colon U_{2}\rightarrow\varphi_{2}(U_{2}))\in\mathcal{P}_{r_{123}}(G_{2},G_{3})$ such that
@@
For any $p_{3}\in s(\varphi_{2},r_{13})$, we have either $p_{3}\in G_{3}\cap{r_{13}}D$ or $\varphi_{2}(p_{3})\in G_{2}\cap{r_{13}}D\subset G_{2}\cap{r_{123}}D$. Even if $p_{3}\in G_{3}\cap{r_{13}}D$, the inequality $|p_{3}-\varphi_{2}(p_{3})|<d_{2}<r_{2}$ implies that $\varphi_{2}(p_{3})\in G_{2}\cap{r_{123}}D$. Therefore, for any $p_{3}\in s(\varphi_{2},r_{13})$, we obtain $\varphi_{2}(p_{3})\in G_{2}\cap{r_{123}}D$, which allows us to restrict the domain of $\varphi_{2}$ to a smaller neighborhood $U$ that contains $s(\varphi_{2},r_{13})$ and on which the composition $\varphi_{1}\circ\varphi_{2}$ is well-defined.
@@
Finally, we prove that the topology induced by $\{d_{r}\}_{r>0}$ coincides with the $C^{0}$-topology defined in [MR2784914], Definition 2.5. First, we show that for any $G\in\Phi{(\mathbb{R}^{N})}$, any compact set $K\subset\mathbb{R}^{N}$, and any $\varepsilon>0$, there exist $r^{\prime}>0$ and $\varepsilon^{\prime}>0$ such that